# Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned

### v1

@@ 6.1 Measurement of Effectiveness in Production
Accurately measuring the effectiveness of hallucination detection and mitigation and the values bringing to customers in a production environment is very challenging. We have designed a system which applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content. However, the GPT4-based measurement has limitation and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
@@ 6.2 Handling Multilingual and Long Source Documents
We have add several major non-English training data in NLI and SBD models to support the multilingual use case and using segmentation based approach to deal with long source documents. However, managing multilingual even cross-lingual inputs and long source documents adds another layer of complexity. Also it is still an open research problems how to do effective modeling of the long source document.
6.3 Deep Customization Needs for Various Scenarios
Different usages and user circumstances call for tailored adjustments. In the production, we’ve calibrated our ensemble-based detection for greater accuracy with a reduced block rate, while also adjusting our rewriting for decreased latency at the cost of some mitigation power. As we move forward, incorporating additional domain or task-specific datasets may enhance our detection capabilities. Concurrently, researching novel detection techniques or focusing exclusively on detecting certain types or severe cases of hallucinations might be more practical for real-world applications. Furthermore, we should consider investigating more efficient mitigation strategies involving smaller language models.
@@ 7 Conclusions
To conclude, we developed a reliable and high-speed production service to detect and mitigate intrinsic or close-world hallucinations in LLMs. Our service combines named entity recognition (NER), natural language inference (NLI), span-based detection (SBD) models, and a decision tree analysis to identify diverse set of hallucinations with high accuracy. We also develop an efficient rewriting mechanism to ensure the quality and fidelity of LLM outputs.
Hallucination is a key problem facing by LLM applications. Our journey from idea to deployment revealed several challenges in delivering high-quality solutions under real-world constraints. We had to balance latency, availability, and quality in production settings.
Our work has broad applicability and potential for future impact, as LLMs are used in diverse sectors that require robust hallucination verification mechanisms. Our service can adapt to different LLM applications, each with different challenges and demands. A key next step is to address extrinsic or open-world hallucinations, which require integrating external knowledge bases and information verification systems.

### v2

@@ 6.1 Measurement of Effectiveness in Production
Accurately measuring the effectiveness of hallucination detection and mitigation, as well as the value they bring to customers in a production environment, is very challenging. We have designed a system that applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content. However, the GPT4-based measurement has limitations and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
@@ 6.2 Handling Multilingual and Long Source Documents
We have incorporated major non-English training datasets into our NLI and SBD models to support multilingual use cases and are utilizing a segmentation-based approach to manage long source documents. However, handling inputs and outputs in different languages and their extensive combinations remains challenging. Additionally, developing effective models for processing long source documents continues to be an open research problem, requiring further exploration and innovation.
6.3 Deep Customization Needs for Hallucination Handling
Different user circumstances call for tailored adjustments to hallucination handling. For example, to meet the production needs, we’ve calibrated our ensemble-based detection for greater precision with a reduced block rate (or trigger rate) to avoid the availability issue, while also adjusting our rewriting for decreased latency at the cost of some mitigation power. However, there might be another setting, where we need a different balance of quality, latency and cost. Also, adapting to special domain or task or handling a specific types of hallucinations can also be great directions to explore.
@@ 7 Conclusions
In this paper, we introduce a novel framework that can detect and mitigate intrinsic hallucinations, characterized by outputs not supported by grounding documents in LLMs. Our detection approach leverages the combined strengths of NER, NLI, and novel sequence labelling (SBD), and Decision Tree to detect as much as hallucination as possible. We further developed an effective LLM-based mitigation solution that blance the quality and latency.
We detail the core elements of our framework and underscore the paramount challenges tied to response time, availability, and performance metrics, which are crucial for real-world deployment of these technologies. Our extensive evaluation, utilizing offline data and live production traffic, confirms the efficacy of our proposed framework and service.